OpenStudy (anonymous):

Please help! I'll give you a medal<3 1. Which ordered pair is not a solution of y = –12x – 3? A. (-3,33) B. (8, -98) C. (-9, 105) D) (3, -39)

OpenStudy (tkhunny):

Have you considered substituting the candidates? Candidate: (-3,33) Substitute: 33 = -12(-3) + 3 = 36 + 3 = 39 <== Whoops! That's no good. Throw out the candidate (-3,33). You do the rest.

OpenStudy (anonymous):

the first one works....

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

tk check your math

OpenStudy (tkhunny):

Check my Transcription!! Substitute: 33 = -12(-3) - 3 = 36 - 3 = 33 <== Woo-hoo! That one works!

OpenStudy (anonymous):

now you got it!